OpenStudy (anonymous):

If the graph of the equation ax + 3y = 9 is perpendicular to the graph of the equation 3x + y = -4, find the value of a.

OpenStudy (anonymous):

When you multiply the slopes of two linear equations, and you get a -1, then the lines are perpendicular. So rearrange your first equation, so that it is in this form: y=mx+b m is the slope of the equation. So you you would have this: -a/3x +3 = y So -a/3 is the slope in this equation. Rearrange the second equation to the same form. So it looks like this: -3x-4= y In this equation, the slope is -3. So now you set the product of the two slopes equal to -1 So it looks like this: (-3)(-a/3)= -1 Now solve for a. -a/3=1/3 a=1